OpenStudy (anonymous):

Differentiate (X5-5X+8)(x3+7x+9) in three ways mentioned below (i) By using product rule. (ii) By expanding the product to obtain a single polynomial. (iii By logarithmic differentiation. Do they all give the same answer?

OpenStudy (mimi_x3):

Is it: \[\large(x^5-5x+8)(x^3+7x+9)\]?

OpenStudy (anonymous):

yes,it is..help me now

OpenStudy (mimi_x3):

Well, have you tried the product rule?

OpenStudy (mimi_x3):

u = x^5 -5x +8 => u'=5x^4-5 v = x^3+ 7x+9 => v' = 3x^2+7 then product rule..

OpenStudy (anonymous):

how by logarithmic diffrenciation?